# Directional Sign Loss: A Topology-Preserving Loss Function that Approximates the Sign of Finite Differences

### v1

@@ 3.3 Demonstration
(a) Original vs reconstructions values.
(b) Original vs reconstructions cumulative signs.
Figure 2: Example originals and reconstructions of the generated sinusoidal wave dataset. The cumulative signs (2(b)) increase by 1 when the sign of the finite difference is positive, and decreases if it is negative. The combination of training with 1 MSE + 128 DSL results in reconstructions that have similar shape to the original and have movement directions like those of the original.
@@
To illustrate DSL’s behavior, we use a simple example. We trained a simple autoencoder consisting of 2-layer perceptrons with $\mathrm{ReLU}$ between the layers for both the encoder and the decoder on a generated dataset. The dataset consists of 2048 points on one cycle of a sinusoidal wave, starting at some random point in the cycle, i.e. $\sin\left([0:2\pi]+U(0,2\pi)\right)$, scaled by $e^{[-1:1]}$ with 0.5 probability and $e^{[1:-1]}$ with 0.5 probability. The autoencoder compressed the data to a latent dimensionality of 2.
@@
Figure 2(a) shows an example and its reconstruction after training with the conventional MSE loss, DSL only with $s=32$, and a combination of 1 MSE + 128 DSL. This may appear imbalanced, but was necessary for DSL to have an obvious effect on training, as the MSE loss value was usually about 150 times the DSL value (Table 1). As expected, using MSE loss gives the closest reconstruction, but exhibits many direction changes compared to the original. Using only DSL, one may expect it to ideally go up or down linearly until the crest or trough of the wave, but this is not what happens. The reconstruction stays close to 0 while following the direction of the original curve, occasionally jumping back. This tendency arises from a combination of the low latent dimensionality and the properties of DSL. Outputs are encouraged to stay near 0 to avoid large penalties from mismatches and to allow the parameters to easily change the direction of the trend, as all it has to do is get the direction right. While the jump across 0 may appear large from the perspective of the absolute value of finite differences, they only cost one wrong sign in DSL.
@@
Figure 2(b) shows the cumulative signs, which increase by 1 when the sign of the finite difference is positive, and decrease if it is negative. This can be thought of as the ideal output using only DSL. We expect constantly increasing or decreasing cumulative signs, which change directions at a sharp point, i.e. a sawtooth wave which is reflected in the original. Using the combination of the loss functions, the reconstruction exhibits characteristics of both previous cases, showing a curve that resembles the shape of the original curve, with mostly the same trends, and a more sawtooth-like cumulative signs curve.
Training Loss Function Evaluation Metrics After Training
MSE DSL Exact DSL Persistence
MSE 23.3221 0.5920 968.1504 0.6992
DSL 1887.9517 0.1471 194.4619 2.0775
1 MSE + 128 DSL 185.3973 0.0944 429.5840 0.4317
Table 1: Loss-like metrics after training the 2-layer perceptron autoencoder on the sinusoidal dataset with different loss functions. The first column lists the loss function used for training, while the remaining columns contain evaluation metrics after training. “Persistence” denotes the Wasserstein distance between persistence diagrams of the inputs and reconstructions. The reconstructions from the model trained with 1 MSE + 128 DSL had the closest persistence diagrams to those of the originals.
Table 1 shows several metrics after training with different loss functions. In the table, “Exact DSL” refers to the non-differentiable error metric mentioned in § 3.1 that uses the $\mathrm{sign}$ function. As expected, the model trained with only DSL performed best on this metric. “Persistence” refers to the Wasserstein distance between persistence diagrams of the inputs and reconstructions, a common way to compare topological features. We only created 0-dimensional persistence diagrams (not to be confused with tensor dimensionality), as higher-dimensional diagrams were impractically expensive (Appendix 5.1.1). The model trained on only DSL, although encouraged to preserve trends, did not perform best in persistence diagram distance because persistence diagrams account for codomain values of critical points, and are thus sensitive to raw values. The model trained on the 1 MSE + 128 DSL mixture performed best on this metric because it balanced preserving values and preserving trends.
@@
This demonstration shows two extremes in a continuum of loss function choice. From the perspective of conventional distances, MSE does well and DSL does poorly, while from the perspective of behavior comparison, DSL does well while MSE does poorly. Note that we used a low latent dimensionality of 2 to emphasize the differences between loss functions. If it were increased to 4, all reconstructions except that using DSL only would closely overlap and follow the original. This suggests that in the typical case, where the preservation of trends is relevant, a mixture of loss functions would give the most desirable result.
@@ 4 Experiments
DSL can be used with MSE loss in most learnable compressor training processes. To assess its impact, we trained autoencoders on 1-dimensional stock prices, 2-dimensional images, and 3-dimensional physics simulation states. Hyperparameters are in Table 4. For details, see the code.
@@ 4.1 Stock prices
Stock prices are challenging to compress because price changes resemble Brownian motion (Samuelson, 2015). Certain trading strategies focus on the direction of price changes instead of the magnitude or absolute price (Bakhach et al., 2016; Aloud, 2015). We downloaded minute-level prices of NVDA from 2024-12-24 to 2025-01-22, totaling 15360 prices. We sampled subsequences of 64 minutes, normalized them to start at 0, and divided them by the overall standard deviation. The autoencoder had a 2-layer perceptron encoder and decoder with $\mathrm{ReLU}$ between layers. The latent dimensionality was 32.
(a) Original normalized prices vs reconstructions values.
(b) Original vs reconstructions cumulative signs.
Figure 3: Original normalized stock prices vs reconstructions after training a simple autoencoder using MSE, DSL, or a combination of 1 MSE + 100 DSL. In 3(b), “agreement” is the proportion of finite difference signs that are the same as those of the original. While in 3(a), the MSE and combination curves both follow the original closely, in 3(b), the DSL curve moves in the same direction most often, with the combination in second place.
Training Loss Function Evaluation Metrics After Training
MSE DSL Exact DSL Persistence
MSE 3.3800 0.5847 21.4571 0.7007
DSL 177.4241 0.5410 20.0137 2.3836
1 MSE + 100 DSL 4.3236 0.5677 20.9642 0.8034
Table 2: Loss-like metrics after training the 2-layer perceptron autoencoder on NVDA stock prices. Adding DSL to the loss function leads to better performance on the exact DSL metric, but results in greater persistence diagram distance, in which the slightly higher proportion of correct directions exhibited by the DSL and mixture models was not enough to compensate for the deviation from the raw prices.
Figure 3 shows an example reconstruction after training with MSE loss only, DSL only with $s=16$, and a combination of 1 MSE + 100 DSL. As expected, the reconstruction using only DSL stayed near 0, but the other reconstructions followed the original price closely. The differences between them are more obvious in Figure 3(b), which illustrates that the DSL reconstruction more frequently moved in the same direction as the original compared to the others. The combination loss reconstruction showed a lower agreement rate than the DSL reconstruction but was still higher than when using MSE loss alone.
Table 2 shows loss-like metrics after training. As we expect, the models trained using MSE and DSL perform best on the MSE and (exact) DSL metrics respectively. What may be unexpected is that persistence diagram distance was best when using MSE alone. This can be explained by considering the data: the Brownian-like motion of stock prices appears as unstructured noise on the relatively short-term scale of minutes, leading to a high proportion of incorrect directions in all reconstructions. The MSE reconstructions, and to some extent the 1 MSE + 100 DSL reconstructions, made up for it by having reconstructions with closer values to the original.
@@ 4.2 Images
The field of learnable data compressors has had the most study in the domain of image compression, particularly using variational autoencoders (VAEs), which are autoencoders whose latent representations are encouraged to be normally distributed. This allows generating new images by sampling from the normal latent distribution. We trained a quantization-aware ResNet-based VAE designed for image compression (Duan et al., 2024), using MSE loss, DSL, and 0.5 MSE + 0.5 DSL. We used Algorithm 2 on COCO 2017 (Lin et al., 2014), the training dataset, to find a sharpness of 9.6.
Table 3 shows loss-like metrics after training using MSE, DSL, or 0.5 MSE + 0.5 DSL. The table introduces the image-oriented metrics bits per pixel (BPP) and peak signal-to-noise ratio (PSNR). BPP is the average number of bits needed to encode each pixel, computed from the negative log-likelihood of the image under the model and the Kullback–Leibler (KL) divergence between the approximate posterior and the prior. PSNR is the scaled logarithm of the squared dynamic range divided by the MSE, a measurement of the distortion or noise in a signal compared to a reference, often used in image and audio processing. The model trained using 0.5 MSE + 0.5 DSL achieved the best scores on all metrics except BPP. This is likely due to the training scheme which attempts to balance BPP and loss by randomly choosing a reconstruction loss factor per training iteration, relative to a constant KL divergence loss factor (Duan et al., 2024).
Training Loss Function Evaluation Metrics After Training
BPP MSE DSL Exact DSL Persistence PSNR
MSE 0.3553 0.0012 0.9953 1197730.9583 0.1933 29.5401
DSL 2.4361 0.0369 0.9741 1124641.7917 0.4650 14.9261
0.5 MSE + 0.5 DSL 7.2529 0.0001 0.1854 456662.6250 0.1088 41.5673
Table 3: Evaluation metrics after training the model with different loss functions. Bits per pixel (BPP) quantifies the compression and quality trade-off in a VAE, and peak signal-to-noise ratio (PSNR) is a measurement of the distortion or noise in a signal compared to a reference, where higher is better. With the combination 0.5 MSE + 0.5 DSL, the model achieves the best performance on all metrics except BPP.
Figure 4 shows an example image from the Kodak Lossless True Color Image Suite (Eastman Kodak Company, 1999) and its reconstructions, using the lowest reconstruction loss weight of 16, encouraging minimum bit rate. The only visible differences among the reconstructions are that using only DSL, the reconstruction tends to stay close to the middle of the dynamic range, resulting in a faded-looking reconstruction. This reflects the tendency of the reconstruction to stay near zero in § 3.3.
(a) Original
(b) MSE only
(c) DSL only
(d) 0.5 MSE + 0.5 DSL
Figure 4: An example image and its reconstructions from the validation dataset (Eastman Kodak Company, 1999). Using only DSL (4(c)), the reconstruction tends to stay close to the middle of the dynamic range, resulting in a faded-looking reconstruction. We cannot see any differences between the others and the original.
@@ 4.3 Physical simulations
Shallow water simulations model fluid flow under the assumption that the water surface dimensions are much larger than its depth, which allows the simulations to avoid modeling vertical fluid velocity. These simulations solve a set of partial differential equations that describe the conservation of mass and momentum in a fluid layer, and can output surface height and fluid velocities on horizontal axes. We used a simulation to generate data for evaluating DSL on a 3-dimensional dataset. Alternative datasets such as SDRBench (Zhao et al., 2020) had only hundreds of examples or were best described as 2-dimensional with multiple channels (e.g. images). Our simulation initializes a shallow container of water with a Gaussian bump-shaped high spot in a random location. As time steps progress, the bump falls and the water moves around. We take snapshots of the water surface height and return batches of examples shaped (number of snapshots, container length, container width). For this experiment, examples are shaped (8, 64, 64). On this dataset, we trained an autoencoder with two 3-dimensional convolutional layers for both the encoder and decoder, with $\mathrm{ReLU}$ between layers. DSL sharpness was 16.
Figure 5: Loss-like metrics after training the 3-dimensional autoencoder on shallow water simulations using a mixture of MSE and DSL, with ranging from 1.0 MSE + 0.0 DSL to 0.0 MSE + 1.0 DSL in steps of 0.1. Because the metrics use different units, some are scaled so that their trends are easily comparable. Persistence diagram distance is minimized when using only DSL, but MSE is worst using that scheme. 0.9 MSE + 0.1 DSL (near the left side) appears to balance all metrics while occupying the local minimum in exact DSL.
Figure 5 shows loss-like metrics after training on mixtures of MSE and DSL at varying proportions. Persistence diagram distance appears to peak at 0.7 MSE + 0.3 DSL. This can be interpreted as not achieving reconstruction fidelity in terms of raw values while also not faithfully reproducing the directional trends. At other points in the domain, we observe expected training behavior, where MSE loss is minimized when MSE weight is highest, and DSL loss and similar metrics are minimized where DSL weight is highest.
(a) MSE only
0.5 MSE + 0.5 DSL
(c) DSL only
Figure 6: Reconstructions using some mixtures of MSE and DSL. For the full figure, see Appendix 7. Block-shaped patterns appear in the reconstructions using a high proportion of DSL, corresponding to the upsampling factor of 2 in the two convolutional decoders.
@@
The minimum persistence diagram distance is achieved when using only DSL, but this results in the jumping-across-zero patterns that we observed in § 3.3, as visible in Figure 6(c). These patterns, which repeat every 4 elements, correspond to the mapping between the convolutional decoders and their inputs, where each input element corresponds to a $4\times 4\times 4$ block in the reconstruction. This suggests that the convolutional decoder layers have learned to output the correct water surface height gradients while the other layers may have focused on other features.
5 Conclusion and Future work
We introduced a novel loss function, directional sign loss (DSL) and investigated its properties and effects on training a variety of autoencoders on diverse datasets. We found that DSL, when used in combination with MSE loss, can serve as a proxy for common non-differentiable and computationally expensive topological feature descriptors, namely the Wasserstein distance between persistence diagrams, and can encourage autoencoders to preserve topological features such as critical points in their reconstructions while maintaining or minimally sacrificing performance on traditional metrics.
@@ 5 Conclusion and Future work
For future work, because we have observed that learning behavior is sensitive to the balance of DSL and traditional loss functions, we anticipate a way to find an optimal balance, facilitating hyperparameter search. Also, although adjusting the per-dimensional weights of DSL can enable its use for vector fields and higher-dimensional fields, we have not explored this aspect. Finally, while DSL can be introduced into the training process of any typical autoencoder, we have not yet explored model architectures besides MLPs and CNNs. A task requiring a transformer-based autoencoder would bring additional insights into DSL’s effects.
@@ 5.1.1 Persistence diagram
      If $x_{i}$ is a local minimum (i.e., all its neighbors in the grid have higher values), a new connected component is born. The definition of neighbors for each grid point is its orthogonally and diagonally adjacent elements. For a $d$-dimensional grid, each point has $13^{d}-$ 1 neighbors.
@@ 5.1.2 Example persistence diagram
    Flatten and sort the values: $[1,1,2,2,3,3,4,5,6]$, keeping track of their locations.
@@
The persistence diagram will have points $(1,2),(1,3),(1,4),(1,5),(1,6)$.

### v3

@@ 3.3 Demonstration
To illustrate DSL’s behavior, we use a simple example. We trained 2-layer perceptrons with $\mathrm{ReLU}$ between the layers for both the encoder and the decoder on a generated dataset. The dataset consists of 2048 points sampled from one cycle of a sinusoidal wave, starting at a random phase in the cycle, i.e., $\sin([0,2\pi]+U(0,2\pi))$. With probability 0.5, this waveform is multiplied by an exponential function that smoothly increases from $e^{-1}$ to $e^{1}$; otherwise, it is multiplied by an exponential function that decreases from $e^{1}$ to $e^{-1}$. The autoencoder compressed the data to a low latent dimensionality of 2 (the number of independent features) to emphasize the differences between loss functions.
@@
Figure 3(a) shows an example and its reconstruction after training with the conventional MSE loss, DSL only with $s=32$, and a combination of 1 MSE + 128 DSL. This may appear imbalanced, but was necessary for DSL to have an obvious effect on training, as the MSE loss value was usually about 150 times the DSL value. As expected, using MSE loss gives the closest reconstruction, but exhibits many direction changes compared to the original. Using only DSL, one may expect it to ideally go up or down linearly until the crest or trough of the wave, but this is not what happens. The reconstruction stays close to 0 while following the direction of the original curve, occasionally jumping back. This tendency arises from a combination of the low latent dimensionality and the properties of DSL. Outputs are encouraged to stay near 0 to avoid large penalties from mismatches and to allow the parameters to easily change the direction of the trend, as all it has to do is get the direction right. While the jump across 0 may appear large from the perspective of the absolute value of finite differences, they only cost one wrong sign in DSL.
@@
Figure 3(b) shows the cumulative signs, which increase by 1 when the sign of the finite difference is positive, and decrease if it is negative. This can be thought of as the ideal output using only DSL. We expect constantly increasing or decreasing cumulative signs, which change directions at a sharp point, shown in the original. Using the combination of the loss functions, the reconstruction exhibits characteristics of both previous cases, showing a curve that resembles the shape of the original curve, with mostly the same trends, and a more sawtooth-like cumulative signs curve.
(a) values
(b) cumulative signs
Figure 3: Example originals and reconstructions of the generated sinusoidal wave dataset. The cumulative signs (3(b)) increase by 1 when the sign of the finite difference is positive, and decrease when it is negative. The combination of training with 1 MSE + 128 DSL results in reconstructions that have similar shape to the original and share movement directions.
@@
This demonstration shows two extremes in a continuum of loss function choice. From the perspective of conventional distances, MSE does well and DSL does poorly, while from the perspective of behavior comparison, DSL does well while MSE does poorly. This suggests that in the typical case, where the preservation of trends is relevant, a mixture of loss functions would give the most desirable result. Of course, with sufficient resources (parameters, precision, etc.) either loss function encourages perfect reproduction of critical points. However, with limited resources, we must select different loss functions to emphasize different objectives.
@@ 4 Experiments
DSL can be used with MSE loss in most learnable compressor training processes. To assess its impact, we trained autoencoders on 1-dimensional arrays of stock prices, 2-dimensional images, and 3-dimensional physics simulation states. Key hyperparameters are in Appendix Table 1.
@@ 4.1 Stock prices
Stock prices are challenging to compress because price changes resemble Brownian motion (Samuelson, 2015). Certain trading strategies focus on the direction of price changes instead of the magnitude or absolute price (Bakhach et al., 2016; Aloud, 2015). We downloaded minute-level prices of NVDA from 2024-12-24 to 2025-01-22, totaling 15360 prices. We sampled subsequences of 64 minutes, normalized them to start at 0, and divided them by the overall standard deviation. We trained an MLP on this data with a latent dimensionality of 16.
Figure 4 shows an example reconstruction after training with different combinations of MSE and DSL. As expected, low latent dimensionality (given the Brownian-like data) forces the MSE-dominant models to smooth over the minute-level movements. Figure 5 shows three topological dissimilarity metrics after training: persistence diagram distance, functional distortion distance between merge trees, and pairwise linear correlation distance. When we use persistence diagram distance, we take the average of the order-2 Wasserstein distances among all homology orders up to $d-1$.
(a) normalized prices
(b) cumulative signs
(c) color bar
Figure 4: Original normalized stock prices vs reconstructions after training a simple autoencoder using MSE, DSL, or a combination thereof. While the MSE-dominant models follow the original closest, they smooth over the minute-level movements. The proportions of movement signs that are the same between the original and reconstructions are shown on the right side of Figure 4(b). Models trained on mixtures of MSE and DSL are more often correct in direction than those trained on pure MSE or pure DSL.
(a) persistence distance
(b) merge tree distance
(c) correlation distance
Figure 5: Topological dissimilarity metrics after training an MLP on NVDA stock prices. Squares show means while faint dots show individual examples. Error bars show one standard deviation away from the mean. Note the split horizontal axes; the left sides show a finer scale. The Brownian-like movement of minute-level prices is difficult to capture with high compression, although small proportions of DSL $\approx$ 0.08 help capture movement direction.
@@ 4.2 Images
The field of learnable data compressors has had the most study in the domain of image compression, particularly using variational autoencoders (VAEs), which are autoencoders whose latent representations are encouraged to be normally distributed. This allows generating new images by sampling from the normal latent distribution. We trained a quantization-aware ResNet-based VAE designed for image compression Duan et al. (2024), using different mixtures of MSE and DSL. We used Algorithm 2 on COCO 2017 Lin et al. (2014), the training dataset, to find a sharpness of 9.6.
(a) persistence distance
(b) merge tree distance
(c) correlation distance
(d) color bar
Figure 6: Topological metrics on a convolutional VAE trained on COCO 2017. Note the split axis on Figure 6(b). Each group on each reconstruction weight was trained on the same proportion of MSE and DSL, but the points are offset in the plots, and individual examples are omitted for clarity. Means and error bars are shown instead of individual examples. Reconstruction weight is a multiplier on the reconstruction loss as opposed to the KL divergence of the latent distribution from a normal distribution, a VAE optimization objective. Overall, 0.125 DSL proportion performs best.
Figure 6 shows topological metrics, and because we are compressing natural images, Appendix 6.6 shows visual fidelity metrics and reconstruction examples. Images were cropped to shape 3×64×643$\times$ 64$\times$ 643 × 64 × 64 on center when calculating the inefficient merge tree distance and correlation distance metrics.
@@ 4.3 Physical simulations
Shallow water simulations model fluid flow under the assumption that the water surface dimensions are much larger than its depth, which allows the simulations to avoid modeling vertical fluid velocity. These simulations output surface height and fluid velocities on horizontal axes. We used a simulation to generate data for evaluating DSL on a 3-dimensional dataset. Alternative datasets such as SDRBench Zhao et al. (2020) had only hundreds of examples or were best described as 2-dimensional with multiple channels (e.g. images). Our simulation initializes a shallow container of water with a bump in a random location. As time steps progress, the bump falls and the water moves around. We trained an autoencoder with two 3-dimensional convolutional layers for both the encoder and decoder, with $\mathrm{ReLU}$ between layers. Examples were shaped (3 channels, 8 snapshots, 64 length, 64 width). DSL sharpness was 16 and weight in the channel dimension was 0, treating each example as a tensor field with elements of (water height, velocityx, velocityy).
(a) persistence distance
(b) merge tree distance
(c) correlation distance
Figure 7: Topological measures on the 3D autoencoder trained on shallow water simulation states. Because VTK (Schroeder et al., 2006), the library we used for merge tree construction, supports a maximum of 3 dimensions, we evaluated the three channels separately. Models trained on $\approx$ 0.1 DSL perform the best overall.
@@
Figure 7 shows topological metrics after training on mixtures of MSE and DSL. The best overall performance is achieved at $\approx$ 0.1 DSL proportion, suggesting that the most desirable mixture is nearby. Although the minimum correlation distance is achieved when using pure DSL, it results in obvious jumping-across-zero patterns that we observed in § 3.3. These patterns correspond to the mapping between the convolutional decoders and their inputs, where each input element corresponds to a $\times 4\times 4$ 4 block in the reconstruction (See Appendix 11 for visualizations). This suggests that the convolutional decoder layers have learned to output the correct water surface height gradients while the other layers may have focused on other features.
5 Limitations
@@
As demonstrated in the stock prices experiment (§ 4.1), using DSL may not always have a significant benefit. Because learning behavior is sensitive to the balance between the DSL and traditional loss functions, due to their measuring fundamentally different quantities (e.g. squared errors versus counts of sign differences), a method for efficiently finding an optimal balance would facilitate its use. Finally, while DSL can be introduced into the training process of any typical autoencoder, we have not yet explored model architectures besides MLPs and CNNs. A task requiring a transformer-based autoencoder would bring additional insights into DSL’s effects.
6 Conclusion
We introduced directional sign loss (DSL) and investigated its properties and effects on training a variety of autoencoders on diverse datasets. We found that DSL, when used in combination with MSE loss, can serve as a proxy for common non-differentiable and computationally expensive topological feature descriptors, namely the Wasserstein distance between persistence diagrams, the functional distortion distance between merge trees, and the pairwise linear correlation distance, and can encourage autoencoders to preserve topological features such as critical points in their reconstructions.
Acknowledgments and Disclosure of Funding
Supported in part by NSF CISE Award 2217154.
@@ 6.1.1 Persistence diagram
      If $x_{i}$ is a local minimum (i.e., all its neighbors in the grid have higher values), a new connected component is born. The definition of neighbors for each grid point is its orthogonally and diagonally adjacent elements. For a $d$-dimensional grid, each point has $3^{d}-1$ neighbors.
@@ 6.1.2 Example persistence diagram
    Flatten and sort the values: [1,1,2,2,3,3,4,5,6][1,1,2,2,3,3,4,5,6][ 1 , 1 , 2 , 2 , 3 , 3 , 4 , 5 , 6 ], keeping track of their locations.
@@
The persistence diagram will have points (1,2),(1,3),(1,4),(1,5),(1,6)(1,2),(1,3),(1,4),(1,5),(1,6)( 1 , 2 ) , ( 1 , 3 ) , ( 1 , 4 ) , ( 1 , 5 ) , ( 1 , 6 ).
@@ 6.1.3 Wasserstein distance
6.2 Lower time and higher memory complexity than implemented
When calculating DSL, all $\Delta$ arrays for each dimension can be created in constant time, requiring $O(d\prod_{i=1}^{d}n_{i})$ memory to store. Then, all $\Delta$, containing a total of $d\prod_{i=1}^{d}(n_{i}-1)$ elements, are summed. Ignoring the subtraction of 111 for large $n_{i}$, this takes $O(\log_{2}(d\prod_{i=1}^{d}n_{i}))$ time, as elements must be summed two at a time. Taking the logarithm of the product, we have $O(\log_{2}d+\sum_{i=1}^{d}\log_{2}n_{i})$.
6.3 Loss function time and memory experiments with 3D inputs
Figure 8 shows time and memory usage with 3D cube-shaped input arrays.
(a) time 3D
(b) memory 3D
(c) legend
Figure 8: Time and peak memory usage from 3-dimensional array inputs using one AMD Ryzen 7 7800X3D CPU and one NVIDIA RTX 3090 GPU (24GB). Points are not shown if the time per example exceeds 1 second or GPU memory is exceeded. 3D inputs sizes show the number of elements per side of a cube. Persistence diagram distance was the 2-order Wasserstein distance between order-0 (the cheapest) persistence diagrams.